OpenStudy (barinly):

HELP MEDAL >> NEWWWWWW Due to a leak, the amount of water in a pool is decreasing by 5/8 gal every hour. What is the total change in the number of gallons of water in the pool after 3/4 hr? Enter your answer in the box as a simplified fraction.

OpenStudy (barinly):

@jabez177 @Directrix @steve816 @AaronAndyson @Dr-Wells @kimberlyevens

OpenStudy (barinly):

JELPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPP

OpenStudy (festinger):

multiply the 2 values given together.

OpenStudy (barinly):

11/8 @Festinger

OpenStudy (barinly):

is it right... but i thought we have to find... They said it leaks 5/8 every 1 hr.. so?

OpenStudy (festinger):

It isn't right. You added them together, while you should have multiplied them. Why multiply? Say it leaks 1 gallon per hour. In 3 hours it will have leaked \(1*3=3\) gallons. Similarly here, if it leaks \(\frac{5}{8}\) gallons an hours, then in \(\frac{3}{4}\) hour it would have leaked \(\frac{5}{8}\times\frac{3}{4}\) gallons.

OpenStudy (barinly):

So @Festinger the answer is 11/8?????

OpenStudy (festinger):

No, it isn't \(\frac{11}{8}\). \[\frac{3}{4}\times\frac{5}{8}=\frac{5\times3}{8\times4}=\frac{15}{32}\]

OpenStudy (barinly):

so 2 2/15 is the naswer

OpenStudy (barinly):

@Festinger

OpenStudy (festinger):

No, the answer is 15/32. Why did you invert it?

OpenStudy (barinly):

I am suppose to do it into a simplified form?

OpenStudy (barinly):

@Festinger

OpenStudy (barinly):

Enter your answer in the box as a simplified fraction.

OpenStudy (festinger):

15/32 is already in the simplified form. The numerator (15) is smaller than he denominator (32) and 15 and 32 have no other common multiples (except for the number 1).

OpenStudy (barinly):

oh yeah thxs